# Verbalized Algorithms

### v1

@@ 1 Introduction
Imagine receiving thousands of customer complaints and having to quickly sort them by how urgently they need support. Feeding such a large set of input strings blindly into an LLM is ineffective and impractical, both due to its limited context length and its limited ability to perform sorting tasks. LLMs provide no guarantees on the correctness of their outputs for computational reasoning tasks, such as sorting, beyond anecdotal, empirical observations that they sometimes do.
@@
In this work, we propose an abstract framework called “verbalized algorithms”. Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the initial algorithm, thereby largely retaining the theoretical guarantees of the original algorithm.
This has advantages over another common approach to similar tasks that we call the “formalization-based approach”. Leveraging the machine translation ability of LLMs, it converts an informal task representation into a formal, symbolic representation, solves it with a specialized combinatorial solver, and then converts the solution back to the informal, natural language representation for the user. Existing work based on this approach includes LLM-based generation of PDDL (McDermott, 2000) for classical planning (Xie, 2020; Guan et al., 2023; Li et al., 2024; Fine-Morris et al., 2024; Oswald et al., 2024), Prover9 input language for theorem proving (Olausson et al., 2023), or SMT solver input for optimization (Hao et al., 2024). The issue with this approach is twofold: First, it requires grammar-following capability for a complex target formal language. Second, while the downstream solver requires a semantically and logically accurate input, the vagueness and informality in the natural language input often cause LLMs to generate inaccurate and inconsistent translations.
@@ 2 Verbalized Algorithms
While whether LLMs can reason or not is up for both philosophical and scientific debates, it is more widely accepted that they can solve some much simpler tasks with high fidelity, such as text-based classification and question-answering. Assuming those capabilities, we propose verbalized algorithms (VAs), an LLM-based reasoning strategy that limits the scope of the LLM’s usage to simple tasks.
@@
VAs have three categories: Naive VAs, which treat the LLM’s answer as ground truth, Robust VAs, which treat it as a noisy oracle with a certain error rate and quantify the rate of error in the final output, and Probabilistic VAs, which use the probability/logits of LLM outputs to explicitly quantify the uncertainty of the result and return a distribution of outputs. In this workshop paper, we focus on the naive and the robust variants due to space, leaving the probabilistic variants for future work.
2.1 Verbalized Sorting
@@
Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle. Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query such as “Is X better than Y? X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no”. The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively. We evaluated the accuracy of several non-VA and VA approaches. Method, dataset, and prompt details are available in the appendix Sec. A. All results are averages of 5 random seeds.
@@
Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard and Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings. Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list.
@@
Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores. It is a form of formalization-based approach (the list of scores represents the formal problem). We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements.
@@
VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro and Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort. VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput. It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism. VA Robust Powersort/Bitonic Sorting Network makes individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting. For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh and Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963).
@@
First, Table 1 (1) shows how often the constraint-decoding baseline violates the sorting invariance, i.e., how many elements are duplicated or missing in the output on average. While these violations become rarer as the model gets larger, they do not disappear. This highlights the lack of theoretical guarantees in LLM-based reasoning, which does not occur in VAs and formalization-based approaches.
@@
Next, Table 1 (2) shows the average Kendall-Tau scores. Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with a large 32B model, their performance dwindles when the model is smaller. Robust VA with 32B achieved the best results. The strength of VA is most pronounced when using the small 1.7B model, where other approaches failed. Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the literature.
@@
Finally, Table 2 compares the runtime of each approach. VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries). This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature, in contrast to the recent LLM reasoning literature that tends to detach from existing discrete algorithm studies. Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this.
Table 1: (1) Average number of sorting invariance violations by the constraint decoding baseline. “Duplicates” indicates the number of duplicates found in the output, and “missing” indicates the number of input elements that are missing in the output, for 25 input strings. (2) Average Kendall-Tau scores. Top-3 are highlighted in bold. The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
(1) Invariance (2) Accuracy (Kendall-Tau scores) Models Violations Constraint Scoring Naive VA Robust VA ($K=3$) (Qwen3) duplicates missing Decoding i.i.d. AR Bitonic Powersort Bitonic Powersort 1.7B 1.39 9.63 0.00 0.18 0.05 0.30 0.33 0.33 0.44 32B-AWQ 0.49 2.75 0.30 0.57 0.39 0.39 0.56 0.39 0.58
Table 2: Average total elapsed time in seconds for processing all sorting queries. Bitonic sorting network also shows the relative speedup over the sequential sort (Powersort).
Models Scoring Naive VA Robust VA ($K=3$) (Qwen3) Baseline i.i.d. AR Bitonic Powersort Bitonic Powersort 1.7B 7468 630 205 2338 (x2.12) 4975 5060 (x1.45) 7362 32B-AWQ 7280 1430 626 8556 (x2.18) 18679 20350 (x1.30) 26471
2.2 Verbalized Clustering
@@
score $=\textstyle\quad\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
@@
3 Conclusion, Related Work, Discussion and Future Work

### v3

@@ 1 Introduction
Imagine receiving thousands of customer complaints and having to quickly sort them by how urgently they need support. Feeding such a large set of input strings blindly into a large language model (LLM) is ineffective and impractical, both due to its limited context length and its limited ability to perform sorting tasks. LLMs provide no guarantees on the correctness of their outputs for computational reasoning tasks, such as sorting, beyond anecdotal, empirical observations that they sometimes do. Despite the recent improvements, their behavior is fundamentally an approximation of formal reasoning, and thus contains hallucinations.
@@
Such an algorithmic and analytical approach is often missing in LLM-based reasoning literature due to opaqueness of the reasoning mechanism in LLMs. More precisely, it often fails to distinguish the various separate challenges of reasoning tasks, i.e., (1) the ability to accurately perform atomic operations on ambiguous representations, (2) the ability to accurately execute a (not necessarily efficient) algorithm using those operations, (3) the choice of the algorithm to run and its theoretical properties (runtime/memory complexity, parallelism, robustness, soundness, completeness, etc.), and (4) the heuristic policy/hints for a branching algorithm. We argue that most reasoning tasks are informal representations of formal tasks for which an ideal classical algorithm with known complexity guarantees already exists in the literature, thus the effort to further improve (2,3) via additional training is wasteful and unproductive. For example, for the comparison sort, $O(\log n)$ algorithms are asymptotically optimal.
In this work, we propose an abstract framework called “verbalized algorithms”. Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the classical algorithm, thereby largely retaining the original theoretical guarantees.
This has advantages over another common approach to similar tasks that we call the “formalization-based approach”. Leveraging the machine translation ability of LLMs, it converts an informal task representation into a formal, symbolic representation, solves it with a specialized combinatorial solver, and then converts the solution back to the informal, natural language representation for the user. Existing work based on this approach includes LLM-based generation of PDDL (McDermott, 2000) for classical planning (Xie, 2020; Guan et al., 2023; Li et al., 2024; Fine-Morris et al., 2024; Oswald et al., 2024), Prover9 input language for theorem proving (Olausson et al., 2023), or SMT solver input for optimization (Hao et al., 2024). The issue with this approach is threefold: First, it requires grammar-following capability for a complex target formal language. Second, while the downstream solver requires a semantically and logically accurate input, the vagueness and informality in the natural language input often cause LLMs to generate inaccurate and inconsistent translations. Third, the assumptions made in the target formal language limits the scope of the tasks that can be solved by this approach. For example, a LLM-based planning agent that converts the task into PDDL cannot handle an informal task that contains probabilistic, temporal, or partially-observable nature. Verbalized algorithms, on the other hand, can handle them robustly via the general and informal common sense reasoning ability for atomic operations.
@@ 2 Verbalized Algorithms
Auto-regressive generative model is a probability distribution $p({\mathbf{x}})=p({\textnormal{x}}_{1},\ldots,{\textnormal{x}}_{N})$ where the random variable $\mathbf{x}$} represents a vector of integers, typically associated with a word or a sub-word token in natural language. Denoting its subsequence from index $i$ to $j$ as ${\mathbf{x}}_{i:j}$, Transformer-based language models (LMs) Vaswani et al. (2017) decompose it into each conditional distribution modeled by a large set of parameters $\theta$: $p({\mathbf{x}})=p_{\theta}({\mathbf{x}}_{N}|{\mathbf{x}}_{1:N-1})p({\mathbf{x}}_{1:N-1})=\ldots=\prod_{i=j}^{N-1}p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})p({\mathbf{x}}_{1:j})$, where the prefix ${\mathbf{x}}_{1:j}$ is typically called a prompt and each conditional distribution $p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})$ performs a next-token prediction.
To exhibit a trait typically useful for user interaction (e.g. instruction following), LLMs go through a series of parameter-efficient fine-tuning (PEFT) steps after pretraining on a large corpus. For example, in Direct Preference Optimization (Rafailov et al., 2023, DPO), the input is a pair of model outputs in which human users have indicated their preference over another and the model is fine-tuned to follow that preference. For a prompt $x$ and an answer pair $y_{1}$ and $y_{2}$, denote by $y_{1}\succ y_{2}|x$ that $y_{1}$ is preferred over $y_{2}$. DPO learns a Bradley-Terry reward model (Bradley and Terry, 1952, BT-model) for this preference, i.e., $p(y_{1}\succ y_{2}|x)$. The Plackett-Luce reward model (Plackett, 1975; Luce, 1959, PL-model) extends BT to a multinominal case that represents a ranking over choices.
While whether LLMs can reason or not is up for both philosophical and scientific debates, it is more widely accepted that they can solve some much simpler tasks with high fidelity, such as text-based classification and question-answering. This is analogous to say that general BT (and PL) reward models baked in the LLM are sufficient to perform such simple tasks. Assuming those capabilities and thus relying entirely on the accuracy of the reward models, we propose verbalized algorithms (VAs), an LLM-based reasoning strategy that limits the scope of the LLMs to simple tasks.
@@
VAs keep the high-level control flow of $A[\tau]$ intact, while only replacing the oracle $f[\tau]$ with an LLM-based $f[\texttt{str}]$. This preserves the theoretical properties of the classical algorithm $A[\tau]$ if the LLM learned a perfect reward model that is equivalent to the ground truth. If the reward model is inaccurate, a non-classical robust algorithm that assumes noisy oracles can bound the error rate of the output using known theoretical results. VAs can also be combined with other specialized algorithms, such as concurrent algorithms. While our definition is focused on the BT reward model, it naturally extends to PL rewards / categorical oracles.
@@
VAs have three categories: Naive VAs, which treat the LLM’s answer as ground truth, Robust VAs, which treat it as a noisy oracle with a certain error rate and quantify the rate of error in the final output, and Probabilistic VAs, which use the probability/logits of LLM outputs to improve the accuracy, to quantify the uncertainty of the result, or to return a distribution of outputs.
3 Verbalized Sorting
Table 1: Average number of sorting invariance violations by the constraint decoding baseline. “Duplicates” indicates the number of duplicates found in the output, and “missing” indicates the number of input elements that are missing in the output, for 25 input strings.
Models (Qwen3) duplicates missing 1.7B 1.39 9.63 32B-AWQ 0.49 2.75
Table 2: Average Kendall-Tau scores over 5 seeds. Best scores in bold. The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
Models Scoring Naive VA Robust ($K=3$) Symmetrized (Qwen3) Baseline i.i.d. AR Bitonic Powersort Bitonic Powersort Bitonic Powersort 1.7B 0.00 0.18 0.05 0.30 0.33 0.33 0.44 0.33 0.48 32B-AWQ 0.30 0.57 0.39 0.39 0.56 0.39 0.58 0.41 0.60
Table 3: Average total elapsed time in seconds for processing 100 sorting tasks (each with 25 strings). Bitonic sorting network also shows the relative speedup over the sequential Powersort.
Models Scoring Naive VA Robust ($K=3$) (Qwen3) Baseline i.i.d. AR Power. Bitonic Power Bitonic Bitonic+16 1.7B 9279 636 194 5196 2476 (x2.1) 7279 5304 (x1.4) 4728 (x1.5) 32B-AWQ 6798 1406 661 18820 8380 (x2.3) 25302 20461 (x1.2) 17169 (x1.5)
Table 4: Average total elapsed time of naive Powersort in seconds, comparing VLLM without kv-cache sharing, with kv-cache sharing via PagedAttention, and SGLang with RadixAttention.
Models VLLM (PagedAttention) SGLang (Qwen3) disabled enabled (RadixAttention) 1.7B 5196 5403 (x.96) 3812 (x1.4) 32B-AWQ 18820 17421 (x1.1) 14424 (x1.3)
@@ 3 Verbalized Sorting
Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle. Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query such as “Is X better than Y? X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no”. The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively. This task is important in the RAG applications and also called the LLM-based ranking approach. We evaluated the accuracy of several non-VA and VA approaches. Method, dataset, and prompt details are available in the appendix Sec. A. All results are averages of 5 random seeds.
@@
Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard and Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings. Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list. The approach corresponds to listwise ranking approach Ma et al. (2023) in the RAG literature.
@@
Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores. It is a form of formalization-based approach (the list of scores represents the formal problem). We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements. The approach corresponds to pointwise ranking approach Sun et al. (2023); Sachan et al. (2022) in the RAG literature.
@@
VA approaches correspond to pairwise ranking approaches Qin et al. (2024) in the RAG literature, but our contribution focuses on leveraging the theoretical properties of classical algorithms, which has been undermined in the literature. VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro and Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort. VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput. It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism. As an additional optimization, we use the optimal sorting network (Knuth, 1997; Bundala and Závodnỳ, 2014) when the input size is 16 during its recursion. Robust VAs (Powersort/Bitonic) make individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting. For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh and Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963). Finally, Probabilistic VAs (Powersort/Bitonic) instead rely on comparing the probabilities / logits over the yes/no answers to perform each comparison query. While the naive probabilistic VA is equivalent to the naive VA with greedy sampling, we additionally have Symmetrized VA to address two biases in LLM: Positional bias and psycophancy (the tendency to reply “yes”) . Let $p_{12}$ and $p_{21}$ be next-token probabilities for answering “yes” with the original prompt and a prompt whose $x_{1}$ and $x_{2}$ are swapped. Due to the bias, $p_{12}\not=1-p_{21}$. Symmetrized VA tests $\frac{p_{12}+(1-p_{21})}{2}>\frac{(1-p_{12})+p_{21}}{2}$ as a comparison oracle.
Evaluation
@@
First, Table 1 shows how often the constraint-decoding baseline violates the sorting invariance, i.e., how many elements are duplicated or missing in the output on average. While these violations become rarer as the model gets larger, they do not disappear. This highlights the lack of theoretical guarantees in LLM-based reasoning, which does not occur in VAs and formalization-based approaches.
@@
Next, Table 2 shows the average Kendall-Tau scores. Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with the large 32B model, their performance dwindles when the model is smaller. Robust VA with 32B achieved the best results. The strength of VA is most pronounced with the small 1.7B model, where other approaches failed. Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the literature.
@@
Finally, Table 3 compares the runtime of each approach. VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries) or symmetrization (x2 parallel queries). This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature. Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this.
The Effect of Prefix Caching
Each query in VA sort requires a single next-token prediction. In Transformer architecture, it requires $O(t^{2})$ computations for $t$ input tokens. However, State-of-the-Art inference engines such as vllm Kwon et al. (2023) and SGLang Zheng et al. (2024) cache the prefix kv-cache over multiple queries. This is particularly helpful in reducing the inference cost of VA-style queries majority of which share the prefix.
A comparison query in VA sorting consists of a prompt, $x_{1}$, and $x_{2}$. Assume all strings being compared has a fixed length $L=|x_{1}|=|x_{2}|$, and that the length of the fixed prefix is $L_{p}$. In the exhaustive pairwise comparison among $n$ strings ($O(n^{2})$ comparisons, equivalent to insertion sort etc.), a naive inference engine requires $O((L_{p}+2L)^{2})$ runtime per each yes/no query, thus $O(n^{2}(L_{p}+2L)^{2})$ runtime in total, without parallelism.
Assume that an input consists of $N+M$ tokens (prefix and suffix), and the kv-cache of $N$ prefix tokens is already available. Each $N+t$-th new input token attends to N+(t1)N+(t-1) existing tokens, which makes $N+M+1$-th token require $\sum_{t=1}^{M}(N+(t-1))=O(NM+M^{2})$ runtime.
With this in mind, assuming that the fixed prompt already has a kv-cache, the runtime of each comparison query reduces to $O(2L_{p}L+4L^{2})$, totalling $O(L_{p}^{2}+n^{2}(2L_{p}L+4L^{2}))$ runtime. Moreover, since $x_{1}$ is repeated $n$ times over $n^{2}$ comparisons, it further reduces to $O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}((L_{p}+L)L+L^{2}))=O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}(L_{p}L+2L^{2}))$. This is significant because the prompt tends to be much longer than each string.
Table 4 demonstrates the empirical speedup due to kv-cache in two State-of-the-Art inference engines. RadixAttention in SGLang is particularly helpful for VAs whose queries share large prefixes.
4 Verbalized Clustering
@@
score $=\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
@@
5 Related Work
Despite significant recent findings in reasoning models, which combine RL-based self-training with “think tags” (Guo et al., 2025), they rarely explicitly perform systematic backtracking and search.
Whether such systems can reason or not has been hotly debated in the context of LLM-based planning (Valmeekam et al., 2023). Previous work on solving planning tasks using LLM (Liu et al., 2023; Hirsch et al., 2024) demonstrates varying degrees of success and limitations (e.g., failure in mystery due to sensitivity to the vocabulary). Formalization approach to planning is also called a parse-and-solve model (Collins et al., 2022) as opposed to a LLM-as-Planner model that naively generates a plan with an LLM. Thought-of-Search (Katz et al., 2024) simulates a systematic search algorithm, which could be seen as an instance of VAs. To our knowledge, there is no work that simulates traditional graph search algorithms (e.g., A* (Hart et al., 1968), GBFS (Bonet and Geffner, 2001)) in a VA manner, i.e., by replacing all of its atomic operations (e.g., successor generation, heap-based priority queues, hash computation for duplicate detection) with LLM queries, although some work replace some atomic operations, especially successor generations (Katz et al., 2024).
A number of recent papers in the Algorithm with Predictions (Mitzenmacher and Vassilvitskii, 2022, ALPS) community achieved a significant success in augmenting algorithms by data-driven oracles. The particular strength of this line of work is the rigorous theoretical analysis of the error rate. VAs is a form of ALPS specialized in LLM-based oracles, thus combining ideas from ALPS community with VAs is a promising direction for future work.
Combining data-driven heuristics with existing algorithms in the classical AI tasks is also a common practice in the recent literature (e.g., classical planning Núñez-Molina et al. (2024), boolean SAT Duan et al. (2020), logistics tasks Xin et al. (2021)).
6 Conclusion, Discussion and Future Work